# MSHyper: Multi-Scale Hypergraph Transformer for Long-Range Time Series Forecasting

### v1

@@ 4.2.1 Hypergraph Construction
where $\epsilon=(i-1)H_{s}+1$ is the starting node index under the $i\text{th}$ hyperedge of scale $s$ based on original connections. $H_{s}$ is the number of nodes connected by each hyperedge of scale $s$ and $\mathcal{N}(v_{\epsilon}^{s})$ is the neighboring nodes connected to node $v_{\epsilon}^{s}$. Meanwhile, let $\mathcal{G}_{\text{k, ita}}=\{\mathcal{V},\mathcal{E}_{\text{k, ita}}\}$ be the $k$-hop intra-scale hypergraph, where $\mathcal{E}_{\text{k, ita}}=\{\mathcal{E}_{\text{k, ita}}^{s}\}_{s\in\{1,...,S\}}$ contains $k$-hop intra-scale hyperedge sets of different scales. The $i\text{th}$ hyperedge of scale $s$ based on $k$-hop connections $e^{s}_{i,k}\in\mathcal{E}^{s}_{\text{k, ita}}$ is defined as follows:
@@
where $d$ is the starting node index under the $i\text{th}$ hyperedge of scale $s$ based on $k$-hop connections, which can be defined as follows:
@@
where $k$ is the temporal distance between two neighboring nodes. The intra-scale hypergraph $\mathcal{G}_{\text{ita}}=\{\mathcal{V},\mathcal{E}_{\text{ita}}\}$ based on the original intra-scale hypergraph and the $k$-hop intra-scale hypergraph can be formulated as follows:
@@
where $concat$ denotes the concatenation operation. High-order interactions between temporal patterns not only exist between intra-scale temporal patterns, but also between inter-scale temporal patterns, e.g., interactions between hourly and daily patterns, and interactions between daily and weekly patterns. In addition, there are temporal pattern interactions across all scales, e.g., interactions between hourly, daily, and weekly patterns. Therefore, we design the inter-scale hypergraph and mixed-scale hypergraph.
Inter-Scale Hypergraph. As shown in Figure 2(b), the inter-scale hypergraph $\mathcal{G}_{\text{ite}}=\{\mathcal{V},\mathcal{E}_{\text{ite}}\}$ is obtained by concatenating the original inter-scale hypergraph and $k$-hop inter-scale hypergraph. Mathematically, let $\mathcal{G}_{\text{o, ite}}=\{\mathcal{V},\mathcal{E}_{\text{o, ite}}\}$ be the original inter-scale hypergraph, where $\mathcal{E}_{\text{o, ite}}=\{\mathcal{E}_{\text{o, ite}}^{s}\}_{s\in\{1,...,S\}}$ contains original inter-scale hyperedge sets of different scales. The $i\text{th}$ hyperedge of scale $s$ $e^{s}_{i}\in\mathcal{E}^{s}_{\text{o, ite}}$ is defined as follows:
@@
where $l^{s}$ denotes the size of the aggregation window at scale $s$. Then, let $\mathcal{G}_{\text{k, ite}}=\{\mathcal{V},\mathcal{E}_{\text{k, ite}}\}$ be the $k$-hop inter-scale hypergraph, where $\mathcal{E}_{\text{k, ite}}=\{\mathcal{E}_{\text{k, ite}}^{s}\}_{s\in\{1,...,S\}}$ contains $k$-hop inter-scale hyperedge sets of different scales and $e^{s}_{i,k}\in\mathcal{E}^{s}_{\text{k, ite}}$ is defined as follows:
@@
Mixed-Scale Hypergraph. As shown in Figure 2(c), the mixed-scale hypergraph $\mathcal{G}_{\text{mix}}=\{\mathcal{V},\mathcal{E}_{\text{mix}}\}$ is obtained by concatenating the original mixed-scale hypergraph and $k$-hop mixed-scale hypergraph. Mathematically, let $\mathcal{G}_{\text{o, mix}}=\{\mathcal{V},\mathcal{E}_{\text{o, mix}}\}$ be the original mixed-scale hypergraph, where $\mathcal{E}_{\text{o, mix}}$ denotes the original mixed-scale hyperedge set. The $i\text{th}$ hyperedge $e_{i}\in\mathcal{E}_{\text{o, mix}}$ is defined as follows:
@@
where $\delta_{s}=\lceil\epsilon/(1\times\prod_{\alpha=2}^{s}l^{\alpha})\rceil$ is the starting node index of scale $s$ under the $i$th hyperedge based on original connections. Then, let $\mathcal{G}_{\text{k, mix}}=\{\mathcal{V},\mathcal{E}_{\text{k, mix}}\}$ be the $k$-hop mixed-scale hypergraph, where $\mathcal{E}_{\text{k, mix}}$ denotes the $k$-hop mixed-scale hyperedge set. $e_{i,k}\in\mathcal{E}_{\text{k, mix}}$ is defined as follows:
@@
The final hypergraph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ is obtained by considering the abovementioned three hypergraphs, which can be formulated as follows:
@@
4.2.2 Hyperedge Graph Construction
After building the hypergraph, to enhance hypergraph modeling, we build the hyperedge graph to model interactions between hyperedges. Specifically, we build the hyperedge graph by representing hyperedges in $\mathcal{G}$ as nodes and considering the sequential relationship and association relationship. Let $G=(V,E,\mathbf{A})$ be the hyperedge graph, where $V=\{v_{e_{i}}|v_{e_{i}}\in\mathcal{E}\}$ is the node set of $G$ and $E=\{(v_{e_{i}},v_{e_{j}})|v_{e_{i}},v_{e_{j}}\in\mathcal{E}\}$ is the edge set of $G$. $\mathbf{A}$ is the weighted adjacency matrix defined based on the sequential relationship and association relationship between hyperedges.
Sequential Relationship. Within the intra-scale hypergraph, two hyperedges that connect nodes with temporal order exhibit the sequential relationship. For example, as shown in Figure 3(a), hyperedge $e_{1}$ connects the previous three nodes, and hyperedge $e_{2}$ connects the next three nodes. Since long-range time series is a collection of data points arranged in chronological order, intuitively, changes (e.g., increase, decrease, or fluctuation) in the values of nodes connected by $e_{1}$ may influence the values of nodes connected by $e_{2}$. Formally, let $G_{\text{sr}}=(V_{\text{sr}},E_{\text{sr}},\mathbf{A}_{\text{sr}})$ represent the hyperedge graph constructed based on the sequential relationship, where $V_{\text{sr}}=\{v_{{e}_{\text{sr}}}|v_{{e}_{\text{sr}}}\in\mathcal{E}_{\text{ita}}\}\in\mathbb{R}^{{D}_{\text{sr}}\times{D}}$ is the node set and $D_{\text{sr}}$ is the number of intra-scale hyperedges. $E_{\text{sr}}=\{(v_{{e}_{i,\text{sr}}},v_{{e}_{j,\text{sr}}})|v_{{e}_{i,\text{sr}}},v_{{e}_{j,\text{sr}}}\in\mathcal{E}_{\text{ita}},0\leq i-j\leq 1\}$ denotes the edge set constructed based on the sequential relationship. As shown in Figure 3(c), the adjacency matrix constructed based on the sequential relationship $\mathbf{A}_{\text{sr}}\in\mathbb{R}^{D_{\text{sr}}\times D_{\text{sr}}}$ is defined as follows:
@@ 4.2.2 Hyperedge Graph Construction
Association Relationship. Within the inter-scale hypergraph and mixed-scale hypergraph, we define hyperedges with common nodes have an association relationship. As shown in Figure 3(b), $e_{5}$ and $e_{7}$ have an association relationship, as they share a common node $n_{13}$. Formally, let $G_{\text{ar}}=(V_{\text{ar}},E_{\text{ar}},\mathbf{A}_{\text{ar}})$ represent the hyperedge graph constructed based on the association relationship, where $V_{\text{ar}}=\{v_{e_{\text{ar}}}|v_{e_{\text{ar}}}\in\{\mathcal{E}_{\text{ite}},\mathcal{E}_{\text{mix}}\}\}\in\mathbb{R}^{D_{\text{ar}}\times D}$ is the node set of $G_{\text{ar}}$ and $D_{\text{ar}}$ is the number sum of inter-scale hyperedges and mixed-scale hyperedges. $E_{\text{ar}}=\{(v_{{e}_{i,\text{ar}}},v_{{e}_{j,\text{ar}}})|v_{{e}_{i,\text{ar}}},v_{{e}_{j,\text{ar}}}\in\{\mathcal{E}_{\text{ite}},\mathcal{E}_{\text{mix}}\},|v_{{e}_{i,\text{ar}}}\cap v_{{e}_{j,\text{ar}}}|\geq 1\}$ denotes the edge set of $G_{\text{ar}}$. As shown in Figure 3(d), the adjacency matrix based on the association relationship $\mathbf{A}_{\text{ar}}\in\mathbb{R}^{D_{\text{ar}}\times D_{\text{ar}}}$ is defined as follows:
@@
The adjacency matrix $\mathbf{A}$ based on the above two relationships can be defined as follows:
@@
where $\bm{\Gamma}_{1}\in\mathbb{R}^{D_{\text{sr}}\times D_{\text{ar}}}$ and $\bm{\Gamma}_{2}\in\mathbb{R}^{D_{\text{ar}}\times D_{\text{sr}}}$ are matrices consisting entirely of zeros, and $D_{\varphi}$ is the sum of ${D}_{\mathrm{\text{ar}}}$ and ${D}_{\mathrm{\text{sr}}}$.
4.3 Tri-Stage Message Passing Mechanism
@@
Node-Hyperedge Phase. Given the sequences based on the MFE module $\mathbf{X}=\{\mathbf{X}^{1},\mathbf{X}^{2},\ldots,\mathbf{X}^{S}\}\in\mathbb{R}^{N\times D}$, where $N$ denotes the number sum of input time steps and aggregated feature values of different scales. We first get the initialized node embeddings $\bm{\mathcal{V}}=f(\mathbf{X})\in\mathbb{R}^{N\times D}$, where $f$ can be implemented by the multi-layer perceptron (MLP). As shown in Figure 1(c), we get the initialized hyperedge embeddings by the aggregation operation based on the hypergraph $\mathcal{G}$. Specifically, for the $i\text{th}$ hyperedge $e_{i}\in\mathcal{E}$, its initialized embedding is obtained as follows:
$\bm{v}_{e_{i}}=\sum\limits_{v_{j}\in\mathcal{N}(e_{i})}\bm{v}_{j}\in\mathbb{R}^{D},$ (17)
@@
Hyperedge-Hyperedge Phase. After getting initialized hyperedge embeddings, we proceed to update their embeddings through the constructed hyperedge graph. Specifically, for the given initialized hyperedge embeddings $\mathbf{V}=\{\bm{v}_{e_{i}}|\bm{v}_{e_{i}}\in\mathcal{E}\}\in\mathbb{R}^{M\times D}$, we transform it into query $\widetilde{\mathbf{Q}}=\mathbf{VW}^{\text{q}}$, key $\widetilde{\mathbf{K}}=\mathbf{V}\mathbf{W}^{\text{k}}$, and value $\widetilde{\mathbf{V}}=\mathbf{V}\mathbf{W}^{\text{v}}$, where $\mathbf{W}^{\text{q}}$, $\mathbf{W}^{\text{k}}$, and $\mathbf{W}^{\text{v}}$ are learnable weight matrices. For the $i\text{th}$ row $\bm{\widetilde{q}}_{i}\in\widetilde{\mathbf{Q}}$, the updated hyperedge embedding $\widetilde{\bm{v}}_{e_{i}}\in\widetilde{\mathbf{V}}_{\text{e}}$ can be calculated as follows:
$\widetilde{\bm{v}}_{e_{i}}=\sum_{j=1}^{M}\frac{exp(e_{ij})}{\Sigma_{\ell=1}^{M}exp(e_{i\ell})}\widetilde{\bm{v}}_{j}$ (18)
$e_{ij}=\frac{\widetilde{\bm{q}}_{i}\widetilde{\bm{k}}_{j}^{T}}{\sqrt{D}}-(1-\mathbf{A}_{ij})\text{C},$
@@
where $\bm{\widetilde{k}}_{j}^{T}$ denotes the transpose of the $j\text{th}$ row in $\widetilde{\mathbf{K}}$ and $\bm{\widetilde{v}}_{j}$ denotes the $j\text{th}$ row in $\widetilde{\mathbf{V}}$. $\mathbf{A}_{ij}\in\mathbf{A}$ is a binary value and C is a large constant.
@@
Hyperedge-Node Phase. After obtaining the updated hyperedge embeddings, we update the node embeddings by considering all the related hyperedges. Considering the constructed hypergraph $\mathcal{G}$, we use the hypergraph convolution to update the node embeddings. Specifically, the symmetric normalized hypergraph Laplacian convolution can be formulated as follows:
$\bm{\widetilde{\mathcal{V}}}=\sigma(\mathbf{D}_{v}^{-1/2}\mathbf{HD}_{e}^{-1}\mathbf{H}^{T}\mathbf{D}_{v}^{-1/2}\bm{{\mathcal{V}}}\mathbf{{P}})\in\mathbb{R}^{{N}\times{D}_{c}},$ (19)
@@
where $\bm{\widetilde{\mathcal{V}}}$ is the output of the hypergraph convolution and ${D}_{c}$ is the output dimension. $\mathbf{P}\in\mathbb{R}^{D\times D_{c}}$ denotes the learnable parameters and $\sigma$ is the activation function, e.g., LeakyReLU and ELU. To capture the interaction strength of each node $v_{i}\in\bm{\mathcal{V}}$ and its related hyperedges, we dynamically update $\mathbf{H}$ using the node embedding and updated hyperedge embeddings, which can be formulated as follows:
$\mathbf{H}_{ij}^{\text{att}}=\frac{exp(\sigma(f_{t}[\bm{v}_{i},\widetilde{\bm{v}}_{e_{j}}]))}{\sum_{k\in\mathcal{N}_{i}}exp(\sigma(f_{t}[\bm{v}_{i},\widetilde{\bm{v}}_{e_{k}}])))},$ (20)
@@
where [.,.][.,.][ . , . ] denotes the concatenation operation of the node and its related hyperedges. $f_{t}$ is a trainable MLP, and $\mathcal{N}_{i}$ is the neighboring hyperedges connected to $v_{i}$, which can be accessed using the constructed hypergraph $\mathcal{G}$. Then, we use Equation 19 to aggregate pattern information of different scales by replacing $\mathbf{H}$ with $\mathbf{H}^{\text{att}}$. The multi-head attention mechanism is also used to stabilize the training process:
@@
$\bm{\widetilde{\mathcal{V}}}=\bigoplus_{\jmath=1}^{\mathcal{J}}(\sigma(\mathbf{D}_{v}^{-1/2}\mathbf{H}_{\jmath}^{\mathrm{att}}\mathbf{D}_{e}^{-1}\mathbf{H}_{\jmath}^{\mathrm{att}^{T}}\mathbf{D}_{v}^{-1/2}\bm{\mathcal{V}}\mathbf{P}_{\jmath})),$ (21)
@@
where $\bigoplus$ is the aggregation function used for combing the outputs of multi-head, e.g., concatenation or average pooling. $\mathbf{H}^{\text{att}}_{\jmath}$ and $\mathbf{P}_{\jmath}$ are the enriched incidence matrix and the learnable weight matrix of the $j\text{th}$ head, respectively. $\mathcal{J}$ is the number of heads.
4.4 Objective Function

### v2

@@ IV-C1 Hypergraph Construction
where $\epsilon=(i-1)H_{s}+1$ is the starting node index under the $i\text{th}$ hyperedge of scale $s$ based on original connections. $H_{s}$ is the number of nodes connected by each hyperedge of scale $s$ and $\mathcal{N}(v_{\epsilon}^{s})$ is the neighboring nodes connected to node $v_{\epsilon}^{s}$. Meanwhile, we use $\mathcal{G}_{\text{k, ita}}=\{\mathcal{V},\mathcal{E}_{\text{k, ita}}\}$ to represent the $k$-hop intra-scale hypergraph, where $\mathcal{E}_{\text{k, ita}}=\{\mathcal{E}_{\text{k, ita}}^{s}\}_{s\in\{1,...,S\}}$ contains $k$-hop intra-scale hyperedge sets of different scales. The $i\text{th}$ hyperedge of scale $s$ based on $k$-hop connections $e^{s}_{i,k}\in\mathcal{E}^{s}_{\text{k, ita}}$ is defined as follows:
@@
where $d$ is the starting node index under the $i\text{th}$ hyperedge of scale $s$ based on $k$-hop connections, which is defined as follows:
@@
where $k$ is the temporal distance between two neighboring nodes. The intra-scale hypergraph $\mathcal{G}_{\text{ita}}=\{\mathcal{V},\mathcal{E}_{\text{ita}}\}$ based on the original intra-scale hypergraph and the $k$-hop intra-scale hypergraph is defined as follows:
@@
where $concat$ denotes the concatenation operation. High-order interactions between temporal patterns not only exist between intra-scale temporal patterns, but also between inter-scale temporal patterns (e.g., the interactions between hourly and daily patterns, and the interactions between daily and weekly patterns). In addition, there are temporal pattern interactions across all scales (e.g., the interactions between hourly, daily, and weekly patterns). Therefore, we design the inter-scale hypergraph and mixed-scale hypergraph.
Inter-Scale Hypergraph. As shown in Figure 2(b), the inter-scale hypergraph $\mathcal{G}_{\text{ite}}=\{\mathcal{V},\mathcal{E}_{\text{ite}}\}$ is obtained by concatenating the original inter-scale hypergraph and $k$-hop inter-scale hypergraph. We use $\mathcal{G}_{\text{o, ite}}=\{\mathcal{V},\mathcal{E}_{\text{o, ite}}\}$ to represent the original inter-scale hypergraph, where $\mathcal{E}_{\text{o, ite}}=\{\mathcal{E}_{\text{o, ite}}^{s}\}_{s\in\{1,...,S\}}$ contains original inter-scale hyperedge sets of different scales. The $i\text{th}$ hyperedge of scale $s$ $e^{s}_{i}\in\mathcal{E}^{s}_{\text{o, ite}}$ is defined as follows:
@@
where $l^{s}$ denotes the aggregation window size at scale $s$. Then, we use $\mathcal{G}_{\text{k, ite}}=\{\mathcal{V},\mathcal{E}_{\text{k, ite}}\}$ to represent the $k$-hop inter-scale hypergraph, where $\mathcal{E}_{\text{k, ite}}=\{\mathcal{E}_{\text{k, ite}}^{s}\}_{s\in\{1,...,S\}}$ contains $k$-hop inter-scale hyperedge sets of different scales. The $i\text{th}$ hyperedge of scale $s$ based on $k$-hop connections $e^{s}_{i,k}\in\mathcal{E}^{s}_{\text{k,ite}}$ is defined as follows:
@@
The inter-scale hypergraph $\mathcal{G}_{\text{ite}}=\{\mathcal{V},\mathcal{E}_{\text{ite}}\}$ based on the original inter-scale hypergraph and the $k$-hop inter-scale hypergraph is defined as follows:
$\mathcal{G}_{\text{ite}}=concat(\mathcal{G}_{\text{o, ite}},\mathcal{G}_{\text{k, ite}}).$ (11)
Mixed-Scale Hypergraph. As shown in Figure 2(c), the mixed-scale hypergraph $\mathcal{G}_{\text{mix}}=\{\mathcal{V},\mathcal{E}_{\text{mix}}\}$ is obtained by concatenating the original mixed-scale hypergraph and $k$-hop mixed-scale hypergraph. We use $\mathcal{G}_{\text{o, mix}}=\{\mathcal{V},\mathcal{E}_{\text{o, mix}}\}$ to represent the original mixed-scale hypergraph, where $\mathcal{E}_{\text{o, mix}}$ denotes the original mixed-scale hyperedge set. The $i\text{th}$ hyperedge $e_{i}\in\mathcal{E}_{\text{o, mix}}$ is defined as follows:
@@
where $\delta_{s}=\lceil\epsilon/(1\times\prod_{\alpha=2}^{s}l^{\alpha})\rceil$ is the starting node index of scale $s$ under the $i$th hyperedge based on original connections. Then, we use $\mathcal{G}_{\text{k, mix}}=\{\mathcal{V},\mathcal{E}_{\text{k, mix}}\}$ to represent the $k$-hop mixed-scale hypergraph, where $\mathcal{E}_{\text{k, mix}}$ denotes the $k$-hop mixed-scale hyperedge set. The $i\text{th}$ hyperedge based on $k$-hop connections $e_{i,k}\in\mathcal{E}_{\text{k, mix}}$ is defined as follows:
@@
The mixed-scale hypergraph $\mathcal{G}_{\text{mix}}=\{\mathcal{V},\mathcal{E}_{\text{mix}}\}$ based on the original mixed-scale hypergraph and the $k$-hop mixed-scale hypergraph is defined as follows:
$\mathcal{G}_{\text{mix}}=concat(\mathcal{G}_{\text{o,mix}},\mathcal{G}_{\text{k,mix}}).$ (14)
The hypergraph construction module generates the hypergraph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ by considering the intra-scale hypergraph, the inter-scale hypergraph, and the mixed-scale hypergraph, which is defined as follows:
@@
IV-C2 Hyperedge Graph Construction
After building the hypergraph, to enhance hypergraph modeling, we build the hyperedge graph to model the interactions between hyperedges. In GNNs, Chen et al.[10] proposes the concept of the line graph to enhance graph modeling through edge interactions. However, since the line graph is unweighted and considers two edges to be correlated only if the target node of one edge is the source node of the other edge, it lacks in characterizing various edge interaction patterns and is not suitable for modeling hyperedge interactions that connect multiple nodes. Thus, we build the hyperedge graph by representing hyperedges in $\mathcal{G}$ as nodes and considering the sequential relationship and association relationship. We define the hyperedge graph as $G=(V,E,\mathbf{A})$, where $V=\{v_{e_{i}}|v_{e_{i}}\in\mathcal{E}\}$ is the node set of $G$ and $E=\{(v_{e_{i}},v_{e_{j}})|v_{e_{i}},v_{e_{j}}\in\mathcal{E}\}$ is the edge set of $G$. $\mathbf{A}$ is the weighted adjacency matrix defined based on the sequential relationship and association relationship between hyperedges.
Sequential Relationship. Within the intra-scale hypergraph, two hyperedges that connect nodes with temporal order exhibit the sequential relationship. For example, as shown in Figure 3(a), hyperedge $e_{1}$ connects the previous three nodes, and hyperedge $e_{2}$ connects the next three nodes. Since long-range time series is a collection of data points arranged in chronological order, intuitively, changes (e.g., increase, decrease, or fluctuation) in the values of nodes connected by $e_{1}$ may influence the values of nodes connected by $e_{2}$. We use $G_{\text{sr}}=(V_{\text{sr}},E_{\text{sr}},\mathbf{A}_{\text{sr}})$ to represent the hyperedge graph constructed based on the sequential relationship, where $V_{\text{sr}}=\{v_{{e}_{\text{sr}}}|v_{{e}_{\text{sr}}}\in\mathcal{E}_{\text{ita}}\}\in\mathbb{R}^{{D}_{\text{sr}}\times{D}}$ is the node set and $D_{\text{sr}}$ is the number of intra-scale hyperedges. $E_{\text{sr}}=\{(v_{{e}_{i,\text{sr}}},v_{{e}_{j,\text{sr}}})|v_{{e}_{i,\text{sr}}},v_{{e}_{j,\text{sr}}}\in\mathcal{E}_{\text{ita}},0\leq i-j\leq 1\}$ denotes the edge set constructed based on the sequential relationship. As shown in Figure 3(c), the adjacency matrix constructed based on the sequential relationship $\mathbf{A}_{\text{sr}}\in\mathbb{R}^{D_{\text{sr}}\times D_{\text{sr}}}$ is defined as follows:
@@ IV-C2 Hyperedge Graph Construction
Association Relationship. Within the inter-scale hypergraph and mixed-scale hypergraph, hyperedges with common nodes have an association relationship. For example, as shown in Figure 3(b), $e_{5}$ and $e_{7}$ share a common node $n_{13}$. The value changes of nodes connected by $e_{5}$ may influence the values of nodes connected by $e_{7}$ through the common node $n_{13}$. We use $G_{\text{ar}}=(V_{\text{ar}},E_{\text{ar}},\mathbf{A}_{\text{ar}})$ to represent the hyperedge graph constructed based on the association relationship, where $V_{\text{ar}}=\{v_{e_{\text{ar}}}|v_{e_{\text{ar}}}\in\{\mathcal{E}_{\text{ite}},\mathcal{E}_{\text{mix}}\}\}\in\mathbb{R}^{D_{\text{ar}}\times D}$ is the node set of $G_{\text{ar}}$ and $D_{\text{ar}}$ is the number sum of inter-scale hyperedges and mixed-scale hyperedges. $E_{\text{ar}}=\{(v_{{e}_{i,\text{ar}}},v_{{e}_{j,\text{ar}}})|v_{{e}_{i,\text{ar}}},v_{{e}_{j,\text{ar}}}\in\{\mathcal{E}_{\text{ite}},\mathcal{E}_{\text{mix}}\},|v_{{e}_{i,\text{ar}}}\cap v_{{e}_{j,\text{ar}}}|\geq 1\}$ denotes the edge set of $G_{\text{ar}}$. As shown in Figure 3(d), the adjacency matrix based on the association relationship $\mathbf{A}_{\text{ar}}\in\mathbb{R}^{D_{\text{ar}}\times D_{\text{ar}}}$ is defined as follows:
@@
The adjacency matrix $\mathbf{A}$ based on the above two relationships is defined as follows:
@@
where $\boldsymbol{\Gamma}_{1}\in\mathbb{R}^{D_{\text{sr}}\times D_{\text{ar}}}$ and $\boldsymbol{\Gamma}_{2}\in\mathbb{R}^{D_{\text{ar}}\times D_{\text{sr}}}$ are matrices consisting entirely of zeros, and $D_{\varphi}$ is the sum of ${D}_{\mathrm{\text{ar}}}$ and ${D}_{\mathrm{\text{sr}}}$.
IV-D Tri-Stage Message Passing Mechanism
@@
Node-Hyperedge Phase. Given the sequences based on the MFE module $\mathbf{X}=\{\mathbf{X}^{1},\mathbf{X}^{2},\ldots,\mathbf{X}^{S}\}\in\mathbb{R}^{N\times D}$, where $N$ denotes the number sum of input time steps and aggregated feature values of different scales. We first get the initialized node embeddings $\boldsymbol{\mathcal{V}}=f(\mathbf{X})\in\mathbb{R}^{N\times D}$, where $f$ can be implemented by the multi-layer perceptron (MLP). As shown in Figure 1(c), we get the initialized hyperedge embeddings by the aggregation operation based on the hypergraph $\mathcal{G}$. Specifically, for the $i\text{th}$ hyperedge $e_{i}\in\mathcal{E}$, its initialized embedding is defined as follows:
$\boldsymbol{v}_{e_{i}}=\sum\limits_{v_{j}\in\mathcal{N}(e_{i})}\boldsymbol{v}_{j}\in\mathbb{R}^{D},$ (19)
@@
Hyperedge-Hyperedge Phase. After getting initialized hyperedge embeddings, we proceed to update their embeddings through the constructed hyperedge graph. Specifically, for the given initialized hyperedge embeddings $\mathbf{V}=\{\boldsymbol{v}_{e_{i}}|\boldsymbol{v}_{e_{i}}\in\mathcal{E}\}\in\mathbb{R}^{M\times D}$, we transform it into query $\widetilde{\mathbf{Q}}=\mathbf{VW}^{\text{q}}$, key $\widetilde{\mathbf{K}}=\mathbf{V}\mathbf{W}^{\text{k}}$, and value $\widetilde{\mathbf{V}}=\mathbf{V}\mathbf{W}^{\text{v}}$, where $\mathbf{W}^{\text{q}}$, $\mathbf{W}^{\text{k}}$, and $\mathbf{W}^{\text{v}}$ are learnable weight matrices. For the $i\text{th}$ row $\boldsymbol{\widetilde{q}}_{i}\in\widetilde{\mathbf{Q}}$, the updated hyperedge embedding $\widetilde{\boldsymbol{v}}_{e_{i}}\in\widetilde{\mathbf{V}}_{\text{e}}$ is defined as follows:
$\widetilde{\boldsymbol{v}}_{e_{i}}=\sum_{j=1}^{M}\frac{exp(e_{ij})}{\Sigma_{\ell=1}^{M}exp(e_{i\ell})}\widetilde{\boldsymbol{v}}_{j}$ (20)
$e_{ij}=\frac{\widetilde{\boldsymbol{q}}_{i}\widetilde{\boldsymbol{k}}_{j}^{T}}{\sqrt{D}}-(1-\mathbf{A}_{ij})\text{C},$
@@
where $\boldsymbol{\widetilde{k}}_{j}^{T}$ denotes the transpose of the $j\text{th}$ row in $\widetilde{\mathbf{K}}$ and $\boldsymbol{\widetilde{v}}_{j}$ denotes the $j\text{th}$ row in $\widetilde{\mathbf{V}}$. $\mathbf{A}_{ij}\in\mathbf{A}$ is a binary value and C is a large constant.
@@
Hyperedge-Node Phase. After obtaining the updated hyperedge embeddings, we update the node embeddings by considering all the related hyperedges. Considering the constructed hypergraph $\mathcal{G}$, we use the hypergraph convolution to update the node embeddings. Specifically, the symmetric normalized hypergraph Laplacian convolution is defined as follows:
$\boldsymbol{\widetilde{\mathcal{V}}}=\sigma(\mathbf{D}_{v}^{-1/2}\mathbf{HD}_{e}^{-1}\mathbf{H}^{T}\mathbf{D}_{v}^{-1/2}\boldsymbol{{\mathcal{V}}}\mathbf{{P}})\in\mathbb{R}^{{N}\times{D}_{c}},$ (21)
@@
where $\boldsymbol{\widetilde{\mathcal{V}}}$ is the output of the hypergraph convolution and ${D}_{c}$ is the output dimension. $\mathbf{P}\in\mathbb{R}^{D\times D_{c}}$ denotes the learnable parameters and $\sigma$ is the activation function (e.g., LeakyReLU and ELU). To capture the interaction strength of each node $v_{i}\in\boldsymbol{\mathcal{V}}$ and its related hyperedges, we dynamically update $\mathbf{H}$ using the node embedding and updated hyperedge embeddings, which is defined as follows:
$\mathbf{H}_{ij}^{\text{att}}=\frac{exp(\sigma(f_{t}[\boldsymbol{v}_{i},\widetilde{\boldsymbol{v}}_{e_{j}}]))}{\sum_{k\in\mathcal{N}_{i}}exp(\sigma(f_{t}[\boldsymbol{v}_{i},\widetilde{\boldsymbol{v}}_{e_{k}}])))},$ (22)
@@
where [.,.][.,.][ . , . ] denotes the concatenation operation of the node and its related hyperedges. $f_{t}$ is a trainable MLP, and $\mathcal{N}_{i}$ is the neighboring hyperedges connected to $v_{i}$, which can be accessed using the constructed hypergraph $\mathcal{G}$. Then, we use Equation 21 to aggregate pattern information of different scales by replacing $\mathbf{H}$ with $\mathbf{H}^{\text{att}}$. The multi-head attention mechanism is also used to stabilize the training process, which is defined as follows:
@@
$\boldsymbol{\widetilde{\mathcal{V}}}=\bigoplus_{\jmath=1}^{\mathcal{J}}(\sigma(\mathbf{D}_{v}^{-1/2}\mathbf{H}_{\jmath}^{\mathrm{att}}\mathbf{D}_{e}^{-1}\mathbf{H}_{\jmath}^{\mathrm{att}^{T}}\mathbf{D}_{v}^{-1/2}\boldsymbol{\mathcal{V}}\mathbf{P}_{\jmath})),$ (23)
@@
where $\bigoplus$ is the aggregation function used for combing the outputs of multi-head (e.g., concatenation or average pooling). $\mathbf{H}^{\text{att}}_{\jmath}$ and $\mathbf{P}_{\jmath}$ are the enriched incidence matrix and the learnable weight matrix of the $j\text{th}$ head, respectively. $\mathcal{J}$ is the number of heads.
TABLE I: Dataset statistics.
Datasets #Variates ACF values Frequency Information
Flight 7 0.73 Hourly Flight
Weather 21 0.35 10 minutes Meteorological
ETTh1, ETTh2 7 0.46, 0.41 Hourly Temperature
ETTm1, ETTm2 7 0.39, 0.09 15 minutes Temperature
Electricity 321 0.84 Hourly Electricity
Exchange-Rate 8 0.02 Daily Economy
IV-E Objective Function